# Lightweight Transformer for EEG Classification via Balanced Signed Graph Algorithm Unrolling

### v2

@@ 1 Introduction
We study the problem of classifying EEG brain signals from epilepsy patients and healthy subjects. Compared to classical model-based methods, such as k-Nearest Neighbors with dynamic time warping features TASCI2023252 and feature extraction from time–frequency maps 10558582 , deep-learning (DL) models, such as CNN-based 10558582 , 10810442 and Disli2025Epilepsy and more recent transformer-based LIH2023107312 , have achieved state-of-the-art (SOTA) results (e.g., up to the $90\%$ range). However, the transformer model consumes an enormous number of parameters and functions as an uninterpretable black box. Thus, parameter reduction and interpretation of learning models is crucial towards practical implementation on resource-constrained EEG devices.
@@
An alternative paradigm for data learning is algorithm unrolling monga21 : first design an iterative optimization algorithm minimizing a mathematically-defined objective, then “unroll” each iteration into a neural layer, and stack them back-to-back to compose a feed-forward network for data-driven parameter learning. Notably, yu23nips recently unrolls an algorithm minimizing a sparse rate-reduction (SRR) objective into a transformer-like neural net—called “white-box transformer”—that achieves comparable performance as SOTA in image classification, while remaining 100% mathematical interpretable [^1] Common in algorithm unrolling monga21 , “interpretability” here means that each neural layer corresponds to an iteration of an optimization algorithm minimizing a mathematically-defined objective..
@@
Inspired by yu23nips , for the EEG signal classification problem we also build transformers via algorithm unrolling, but from a unique graph signal processing (GSP) perspective ortega18ieee ; cheung18 . GSP studies mathematical tools such as transforms, wavelets, and filters for discrete signals residing on irregular data kernels described by graphs. Recently, Do2024 shows that a graph learning module with edge weight normalization plays the role of self-attention bahdanau14 , and thus unrolling a graph algorithm with graph learning modules inserted yields a transformer-like neural net. However, Do2024 focuses solely on positive graphs that model simple pairwise positive correlations among neighboring pixels in a static image.
@@
For EEG signals, collected samples often exhibit pairwise anti-correlations, which are effectively modeled by negative edges. Though frequencies for general signed graphs (with both positive and negative edges) are not well understood, Dinesh2025 shows that in the special case of balanced signed graphs—with no cycles of odd number of negative edges—frequencies can be rigorously defined: the Laplacian matrix ${\mathcal{L}}^{B}$ of a balanced signed graph ${\mathcal{G}}^{B}$ is a similarity transform of the Laplacian $\mathcal{L}$$^{+}$ of a corresponding positive graph $\mathcal{G}$$^{+}$ (hence they share the same eigenvalues), and the spectra of positive graphs are well understood and utilized in GSP ortega18ieee . Thus, widely studied filters for positive graphs onuki16 ; shuman20 can be readily reused for signals on balanced signed graphs Yokota2025 .
@@
We leverage this fact to build EEG signal denoisers ${\bm{\Psi}}(\cdot)$ as a pretext task [^2] See Appendix A for related works. for later binary classification. Specifically, we first learn a balanced signed graph ${\mathcal{G}}^{B}$ from EEG data; graph balance is ensured during signed edge weight assignment via a novel interpretation of the Cartwright-Harary’s Theorem (CHT) harary53 . Next, we construct an ideal low-pass (LP) filter—parameterized by the lone cutoff frequency $\omega$—for the corresponding positive graph $\mathcal{G}$$^{+}$ to minimize a denoising objective. The ideal LP filter is efficiently implemented via Lanczos approximation susnjara2015 , which we unroll into a filter sub-network. The pair of LP filter / graph learning module is repeated to build a feed-forward network for sparse parameter learning Do2024 ; Cai2025 .
@@
    Extending Do2024 that focuses on positive graphs, we unroll a denoising algorithm for signals on balanced signed graphs with well-defined frequencies—learned directly from data via feature distance learning—into a lightweight and interpretable transformer.
@@
    We implement an ideal LP filter on the positive graph $\mathcal{G}$$^{+}$ corresponding to each learned balanced signed graph ${\mathcal{G}}^{B}$ Dinesh2025 without eigen-decomposition in linear time via Lanczos approximation susnjara2015 , where only the filter cutoff frequency $\omega$ requires tuning from data.
@@
    Compared to SOTA DL methods, we achieve competitive classification performance on EEG signals distinguishing epilepsy patients from healthy subjects, while using significantly fewer parameters (e.g., our scheme achieves $97.6\%$ classification accuracy to transformer-based model LIH2023107312 ’s $85.1\%$, while employing fewer than $1\%$ of the parameters).
@@ 2.1 Graph Signal Processing Definitions
A graph ${\mathcal{G}}({\mathcal{N}},{\mathcal{E}},{\mathbf{W}})$ is defined by a node set ${\mathcal{N}}=\{1,\ldots,N\}$, an edge set $\mathcal{E}$}, and an adjacency matrix ${\mathbf{W}}\in\mathbb{R}^{N\times N}$, where $W_{i,j}=w_{i,j}$ is the weight of edge $(i,j)\in{\mathcal{E}}$ if it exists, and $W_{i,j}=0$ otherwise. In this work, we assume that each edge weight $w_{i,j}$ can be positive or negative to denote positive / negative correlations; $\mathcal{G}$} with both positive and negative edges is a signed graph. We assume also that edges are bidirectional, and thus $w_{i,j}=w_{j,i}$ and $\mathbf{W}$} is symmetric. A combinatorial graph Laplacian is defined as ${\mathbf{L}}\triangleq{\mathbf{D}}-{\mathbf{W}}=\mathrm{diag}({\mathbf{W}}{\mathbf{1}})-{\mathbf{W}}$. To account for self-loops, i.e., $\exists i,W_{i,i}\neq 0$, a generalized graph Laplacian is typically used: ${\mathcal{L}}\triangleq{\mathbf{D}}-{\mathbf{W}}+\mathrm{diag}({\mathbf{W}})$ ortega18ieee . We use these Laplacian definitions for both positive and signed graphs.
@@ 2.2 Graph Laplacian Regularizer
To quantify variation of a signal $\mathbf{x}$} over a graph kernel $\mathcal{G}$}, the graph Laplacian regularizer (GLR) pang17 is commonly used, defined using combinatorial Laplacian $\mathbf{L}$} as
@@ 2.3 Balanced Signed Graphs
A balanced signed graph, denoted by ${\mathcal{G}}^{B}$, is a graph with no cycle of odd number of negative edges. An equivalent definition of graph balance is through node polarities. Each node $i\in{\mathcal{V}}$ is first assigned a polarity $\beta_{i}\in\{1,-1\}$. By the Cartwright-Harary’s Theorem (CHT) harary53 , a signed graph is balanced iff positive/negative edges always connect node-pairs of the same/opposite polarities. In mathematical terms, a signed graph $\mathcal{G}$} is balanced if
@@
Recently, Dinesh2025 proved that there exists a simple similarity transform from the (generalized) graph Laplacian ${\mathcal{L}}^{B}$ of a balanced signed graph ${\mathcal{G}}^{B}$ to a graph Laplacian $\mathcal{L}$$^{+}$ of a corresponding positive graph $\mathcal{G}$$^{+}$, i.e.,
@@
where ${\mathbf{T}}=\mathrm{diag}([-1~~1~~1])$. Given that the graph frequencies of positive graphs are well established [^3] Specifically, eigenvectors of a positive graph Laplacian for increasing eigenvalues have non-decreasing numbers of nodal domains that quantify signal variation across the graph kernel davies00 , and hence can be rightfully interpreted as frequency components (Fourier modes). See Dinesh2025 for details. ortega18ieee , the graph frequencies of balanced signed graphs are also rigorously defined.
@@ 3.1 Balanced Signed Graph Construction
In words, eq.˜11 chooses polarity $\beta_{i}^{*}$ that results in a graph ${\mathcal{G}}^{B}$ more consistent with smooth dataset $\{{\mathbf{x}}^{q}\}_{q=1}^{Q}$, similar in concept as previous works that learn graph Laplacians from assumed smooth signals dong16 ; Kalofolias2016 ; dong19 .
@@
where ${\mathbf{M}}\in\mathbb{R}^{K\times K}$ is a PSD metric matrix, so that $d_{i,j}\geq 0,\forall{\mathbf{f}}_{i},{\mathbf{f}}_{j}$ yang22 .
@@
PSDness: Combinatorial graph Laplacian [^4] A signed graph Laplacian ${\mathbf{L}}^{s}\triangleq{\mathbf{D}}^{s}-\bar{{\mathbf{W}}}^{B}$, where $D^{s}_{i,i}=\sum_{j}|\bar{W}^{B}_{i,j}|$, guaranteed to be PSD can be defined instead dittrich2020 , but a corresponding LP filter would promote negative linear dependence rather than repulsions for negative edges during signal reconstruction. See Appendix C for details. $\bar{{\mathbf{L}}}^{B}=\bar{{\mathbf{D}}}^{B}-\bar{{\mathbf{W}}}^{B}$ may not be PSD due to the presence of negative edges. To ensure PSDness, we leverage the Gershgorin Circle Theorem (GCT) varga04 and add a self-loop of weight $\bar{w}_{i,i}=\delta$ to each node $i$, where $\delta$ is computed as
@@ 3.2 Graph Signal Denoising
Denote by ${\mathcal{S}}_{\omega}({\mathcal{L}}^{+})$ the low-frequency subspace spanned by the first $\omega$ eigenvectors (frequency components) ${\mathbf{V}}_{\omega}=[{\mathbf{v}}_{1};{\mathbf{v}}_{2};\ldots;{\mathbf{v}}_{\omega}]\in\mathbb{R}^{N\times\omega}$ of $\mathcal{L}$$^{+}$ corresponding to the $\omega$ smallest eigenvalues. To denoise observation ${\mathbf{y}}^{+}\in\mathbb{R}^{N}$, we seek a signal ${\mathbf{x}}\in\mathbb{R}^{N}$ in ${\mathcal{S}}_{\omega}({\mathcal{L}}^{+})$ closest to $\mathbf{y}$$^{+}$ in $\ell_{2}$-norm [^5] An alternative is a maximum a posteriori (MAP) denoising formulation using GLR as a signal prior pang17 ; zeng20 ; dinesh20 , i.e., $\min_{\mathbf{x}}\|{\mathbf{y}}^{+}-{\mathbf{x}}\|^{2}_{2}+\mu\,{\mathbf{x}}^{\top}{\mathcal{L}}^{B}{\mathbf{x}}$. However, bai20 shows that the MAP problem—called the E-optimality criterion in optimal design—minimizes the worst-case signal reconstruction, while eq. 18 is the A-optimality criterion that minimizes the average case.:
@@
where $(a)$ is true since columns of $\mathbf{V}$} are orthonormal by the Spectral Theorem hawkins1975 . $g_{\omega}({\mathcal{L}}^{+})={\mathbf{V}}g_{\omega}({\bm{\Lambda}}){\mathbf{V}}^{\top}$ is an ideal LP filter, and $g_{\omega}({\bm{\Lambda}})=\mathrm{diag}([g_{\omega}(\lambda_{1}),\ldots,g_{\omega}(\lambda_{N})])$ has frequency response $g_{\omega}(\lambda_{i})$ defined as
@@
Instead of an ideal LP filter in eq.˜23, we approximate it via Lanczos approximation in complexity ${\mathcal{O}}(N)$ susnjara2015 . In a nutshell, instead of eigen-decomposing a large matrix ${\mathcal{L}}_{m}^{+}\in\mathbb{R}^{N\times N}$, via the Lanczos method we operate on a much smaller tri-diagonal matrix ${\mathbf{H}}_{m}\in\mathbb{R}^{m\times m}$, where $m\ll N$ is the dimension of the approximating Krylov space. We eigen-decompose ${\mathbf{H}}_{m}={\mathbf{Z}}_{m}g({\bm{\Lambda}}_{m}){\mathbf{Z}}_{m}^{\top}$, where $g_{\xi}(\lambda_{i})$ is the approximate LP frequency response for cutoff frequency $\xi=\text{round}(\frac{\omega\,m}{N})$, which we tune from data after unrolling. See Appendix B for details.
@@ 4 Algorithm Unrolling
We implement the graph-based denoising procedure in eq.˜20 and a graph learning module repeatedly; after a solution $\mathbf{x}$$^{*}$ is obtained, representative features $\{{\mathbf{f}}_{i}\}$ are updated (see Section 4.1), resulting in new feature distances $\{d_{i,j}\}$ via eq.˜12, new signed edge weights $\{w_{i,j}\}$ via eq.˜15, and new balanced signed graph Laplacian ${\mathcal{L}}^{B}$ and positive graph Laplacian $\mathcal{L}$$^{+}$ via similarity transform eq.˜3. The concept of iteratively filtering signals, with filter weights updated based on computed signals, is analogous to bilateral filter (BF) in image denoising (tomasi98, ).

### v3

@@ 1 Introduction
We study the problem of classifying EEG brain signals from epilepsy patients and healthy subjects. Compared to classical model-based methods, such as k-Nearest Neighbors with dynamic time warping features Tasci et al. (2023) and feature extraction from time–frequency maps Shen et al. (2024), deep-learning (DL) models, such as CNN-based Shen et al. (2024), Bhandage et al. (2024) and Disli et al. (2025) and more recent transformer-based Lih et al. (2023), have achieved state-of-the-art (SOTA) results (e.g., up to the $90\%$ range). However, the transformer model consumes an enormous number of parameters and functions as an uninterpretable black box. Thus, parameter reduction and interpretation of learning models is crucial towards practical implementation on resource-constrained EEG devices.
@@
An alternative paradigm for data learning is algorithm unrolling Monga et al. (2021): first design an iterative optimization algorithm minimizing a mathematically-defined objective, then “unroll” each iteration into a neural layer, and stack them back-to-back to compose a feed-forward network for data-driven parameter learning. Notably, Yu et al. (2023) recently unrolls an algorithm minimizing a sparse rate-reduction (SRR) objective into a transformer-like neural net—called “white-box transformer”—that achieves comparable performance as SOTA in image classification, while remaining 100% mathematical interpretable [^1] Common in algorithm unrolling Monga et al. (2021), “interpretability” here means that each neural layer corresponds to an iteration of an optimization algorithm minimizing a mathematically-defined objective..
@@
Inspired by Yu et al. (2023), for the EEG signal classification problem we also build transformers via algorithm unrolling, but from a unique graph signal processing (GSP) perspective Ortega et al. (2018); Cheung et al. (2018). GSP studies mathematical tools such as transforms, wavelets, and filters for discrete signals residing on irregular data kernels described by graphs. Recently, Thuc et al. (2024) shows that a graph learning module with edge weight normalization plays the role of self-attention Bahdanau et al. (2014), and thus unrolling a graph algorithm with graph learning modules inserted yields a transformer-like neural net. However, Thuc et al. (2024) focuses solely on positive graphs that model simple pairwise positive correlations among neighboring pixels in a static image.
@@
For EEG signals, collected samples often exhibit pairwise anti-correlations, which are effectively modeled by negative edges. Though frequencies for general signed graphs (with both positive and negative edges) are not well understood, Dinesh et al. (2025) shows that in the special case of balanced signed graphs—with no cycles of odd number of negative edges—frequencies can be rigorously defined: the Laplacian matrix ${\mathcal{L}}^{B}$ of a balanced signed graph ${\mathcal{G}}^{B}$ is a similarity transform of the Laplacian $\mathcal{L}$$^{+}$ of a corresponding positive graph $\mathcal{G}$$^{+}$ (hence they share the same eigenvalues), and the spectra of positive graphs are well understood and utilized in GSP Ortega et al. (2018). Thus, widely studied filters for positive graphs Onuki et al. (2016); Shuman (2020) can be readily reused for signals on balanced signed graphs Yokota et al. (2025).
@@
We leverage this fact to build EEG signal denoisers ${\bm{\Psi}}(\cdot)$ as a pretext task [^2] See Appendix A for related works. for later binary classification. Specifically, we first learn a balanced signed graph ${\mathcal{G}}^{B}$ from EEG data; graph balance is ensured during signed edge weight assignment via a novel interpretation of the Cartwright-Harary’s Theorem (CHT) Harary (1953). Next, we construct an ideal low-pass (LP) filter—parameterized by the lone cutoff frequency $\omega$—for the corresponding positive graph $\mathcal{G}$$^{+}$ to minimize a denoising objective. The ideal LP filter is efficiently implemented via Lanczos approximation Susnjara et al. (2015), which we unroll into a filter sub-network. The pair of LP filter / graph learning module is repeated to build a feed-forward network for sparse parameter learning Thuc et al. (2024); Cai et al. (2025).
@@
    Extending Thuc et al. (2024) that focuses on positive graphs, we unroll a denoising algorithm for signals on balanced signed graphs with well-defined frequencies—learned directly from data via feature distance learning—into a lightweight and interpretable transformer.
@@
    We implement an ideal LP filter on the positive graph $\mathcal{G}$$^{+}$ corresponding to each learned balanced signed graph ${\mathcal{G}}^{B}$ Dinesh et al. (2025) without eigen-decomposition in linear time via Lanczos approximation Susnjara et al. (2015), where only the filter cutoff frequency $\omega$ requires tuning from data.
@@
    Compared to SOTA DL methods, we achieve competitive classification performance on EEG signals distinguishing epilepsy patients from healthy subjects, while using significantly fewer parameters (e.g., our scheme achieves $97.6\%$ classification accuracy to transformer-based model Lih et al. (2023)’s $85.1\%$, while employing fewer than $1\%$ of the parameters).
@@ 2.1 Graph Signal Processing Definitions
A graph ${\mathcal{G}}({\mathcal{N}},{\mathcal{E}},{\mathbf{W}})$ is defined by a node set ${\mathcal{N}}=\{1,\ldots,N\}$, an edge set $\mathcal{E}$}, and an adjacency matrix ${\mathbf{W}}\in\mathbb{R}^{N\times N}$, where $W_{i,j}=w_{i,j}$ is the weight of edge $(i,j)\in{\mathcal{E}}$ if it exists, and $W_{i,j}=0$ otherwise. In this work, we assume that each edge weight $w_{i,j}$ can be positive or negative to denote positive / negative correlations; $\mathcal{G}$} with both positive and negative edges is a signed graph. We assume also that edges are bidirectional, and thus $w_{i,j}=w_{j,i}$ and $\mathbf{W}$} is symmetric. A combinatorial graph Laplacian is defined as ${\mathbf{L}}\triangleq{\mathbf{D}}-{\mathbf{W}}=\mathrm{diag}({\mathbf{W}}{\mathbf{1}})-{\mathbf{W}}$. To account for self-loops, i.e., $\exists i,W_{i,i}\neq 0$, a generalized graph Laplacian is typically used: ${\mathcal{L}}\triangleq{\mathbf{D}}-{\mathbf{W}}+\mathrm{diag}({\mathbf{W}})$ Ortega et al. (2018). We use these Laplacian definitions for both positive and signed graphs.
@@ 2.2 Graph Laplacian Regularizer
To quantify variation of a signal $\mathbf{x}$} over a graph kernel $\mathcal{G}$}, the graph Laplacian regularizer (GLR) Pang and Cheung (2017) is commonly used, defined using combinatorial Laplacian $\mathbf{L}$} as
@@ 2.3 Balanced Signed Graphs
A balanced signed graph, denoted by ${\mathcal{G}}^{B}$, is a graph with no cycle of odd number of negative edges. An equivalent definition of graph balance is through node polarities. Each node $i\in{\mathcal{V}}$ is first assigned a polarity $\beta_{i}\in\{1,-1\}$. By the Cartwright-Harary’s Theorem (CHT) Harary (1953), a signed graph is balanced iff positive/negative edges always connect node-pairs of the same/opposite polarities. In mathematical terms, a signed graph $\mathcal{G}$} is balanced if
@@
Recently, Dinesh et al. (2025) proved that there exists a simple similarity transform from the (generalized) graph Laplacian ${\mathcal{L}}^{B}$ of a balanced signed graph ${\mathcal{G}}^{B}$ to a graph Laplacian $\mathcal{L}$$^{+}$ of a corresponding positive graph $\mathcal{G}$$^{+}$, i.e.,
@@
where ${\mathbf{T}}=\mathrm{diag}([-1~~1~~1])$. Given that the graph frequencies of positive graphs are well established [^3] Specifically, eigenvectors of a positive graph Laplacian for increasing eigenvalues have non-decreasing numbers of nodal domains that quantify signal variation across the graph kernel Davies et al. (2000), and hence can be rightfully interpreted as frequency components (Fourier modes). See Dinesh et al. (2025) for details. Ortega et al. (2018), the graph frequencies of balanced signed graphs are also rigorously defined.
@@ 3.1 Balanced Signed Graph Construction
In words, eq.˜11 chooses polarity $\beta_{i}^{*}$ that results in a graph ${\mathcal{G}}^{B}$ more consistent with smooth dataset $\{{\mathbf{x}}^{q}\}_{q=1}^{Q}$, similar in concept as previous works that learn graph Laplacians from assumed smooth signals Dong et al. (2016); Kalofolias (2016); Dong et al. (2019).
@@
where ${\mathbf{M}}\in\mathbb{R}^{K\times K}$ is a PSD metric matrix, so that $d_{i,j}\geq 0,\forall{\mathbf{f}}_{i},{\mathbf{f}}_{j}$ Yang et al. (2022).
@@
PSDness: Combinatorial graph Laplacian [^4] A signed graph Laplacian ${\mathbf{L}}^{s}\triangleq{\mathbf{D}}^{s}-\bar{{\mathbf{W}}}^{B}$, where $D^{s}_{i,i}=\sum_{j}|\bar{W}^{B}_{i,j}|$, guaranteed to be PSD can be defined instead Dittrich and Matz (2020), but a corresponding LP filter would promote negative linear dependence rather than repulsions for negative edges during signal reconstruction. See Appendix C for details. $\bar{{\mathbf{L}}}^{B}=\bar{{\mathbf{D}}}^{B}-\bar{{\mathbf{W}}}^{B}$ may not be PSD due to the presence of negative edges. To ensure PSDness, we leverage the Gershgorin Circle Theorem (GCT) Varga (2004) and add a self-loop of weight $\bar{w}_{i,i}=\delta$ to each node $i$, where $\delta$ is computed as
@@ 3.2 Graph Signal Denoising
Denote by ${\mathcal{S}}_{\omega}({\mathcal{L}}^{+})$ the low-frequency subspace spanned by the first $\omega$ eigenvectors (frequency components) ${\mathbf{V}}_{\omega}=[{\mathbf{v}}_{1};{\mathbf{v}}_{2};\ldots;{\mathbf{v}}_{\omega}]\in\mathbb{R}^{N\times\omega}$ of $\mathcal{L}$$^{+}$ corresponding to the $\omega$ smallest eigenvalues. To denoise observation ${\mathbf{y}}^{+}\in\mathbb{R}^{N}$, we seek a signal ${\mathbf{x}}\in\mathbb{R}^{N}$ in ${\mathcal{S}}_{\omega}({\mathcal{L}}^{+})$ closest to $\mathbf{y}$$^{+}$ in $\ell_{2}$-norm [^5] An alternative is a maximum a posteriori (MAP) denoising formulation using GLR as a signal prior Pang and Cheung (2017); Zeng et al. (2020); Dinesh et al. (2020), i.e., $\min_{\mathbf{x}}\|{\mathbf{y}}^{+}-{\mathbf{x}}\|^{2}_{2}+\mu\,{\mathbf{x}}^{\top}{\mathcal{L}}^{B}{\mathbf{x}}$. However, Bai et al. (2020) shows that the MAP problem—called the E-optimality criterion in optimal design—minimizes the worst-case signal reconstruction, while eq. 18 is the A-optimality criterion that minimizes the average case.:
@@
where $(a)$ is true since columns of $\mathbf{V}$} are orthonormal by the Spectral Theorem Hawkins (1975). $g_{\omega}({\mathcal{L}}^{+})={\mathbf{V}}g_{\omega}({\bm{\Lambda}}){\mathbf{V}}^{\top}$ is an ideal LP filter, and $g_{\omega}({\bm{\Lambda}})=\mathrm{diag}([g_{\omega}(\lambda_{1}),\ldots,g_{\omega}(\lambda_{N})])$ has frequency response $g_{\omega}(\lambda_{i})$ defined as
@@
Instead of an ideal LP filter in eq.˜23, we approximate it via Lanczos approximation in complexity ${\mathcal{O}}(N)$ Susnjara et al. (2015). In a nutshell, instead of eigen-decomposing a large matrix ${\mathcal{L}}_{m}^{+}\in\mathbb{R}^{N\times N}$, via the Lanczos method we operate on a much smaller tri-diagonal matrix ${\mathbf{H}}_{m}\in\mathbb{R}^{m\times m}$, where $m\ll N$ is the dimension of the approximating Krylov space. We eigen-decompose ${\mathbf{H}}_{m}={\mathbf{Z}}_{m}g({\bm{\Lambda}}_{m}){\mathbf{Z}}_{m}^{\top}$, where $g_{\xi}(\lambda_{i})$ is the approximate LP frequency response for cutoff frequency $\xi=\text{round}(\frac{\omega\,m}{N})$, which we tune from data after unrolling. See Appendix B for details.
@@ 4 Algorithm Unrolling
We implement the graph-based denoising procedure in eq.˜20 and a graph learning module repeatedly; after a solution $\mathbf{x}$$^{*}$ is obtained, representative features $\{{\mathbf{f}}_{i}\}$ are updated (see Section 4.1), resulting in new feature distances $\{d_{i,j}\}$ via eq.˜12, new signed edge weights $\{w_{i,j}\}$ via eq.˜15, and new balanced signed graph Laplacian ${\mathcal{L}}^{B}$ and positive graph Laplacian $\mathcal{L}$$^{+}$ via similarity transform eq.˜3. The concept of iteratively filtering signals, with filter weights updated based on computed signals, is analogous to bilateral filter (BF) in image denoising (Tomasi and Manduchi, 1998).